OpenStudy (anonymous):

Which of the following is a solution for the absolute value inequality |2x-3|>6?

OpenStudy (jdoe0001):

when you have => |2x-3|>6 what's really meant is +1(2x-3)>6 AND -1(2x-3)>6 so solve those 2 and get the values for "x" :)

OpenStudy (anonymous):

i dont really know how to solve this. :/ can you help please?

OpenStudy (jdoe0001):

well, let's try the 1st one how would you solve +1(2x-3) = 6

OpenStudy (jdoe0001):

what would "x" give you?

OpenStudy (anonymous):

Um, do you distribute the 1? Or minus the one on both sides?

OpenStudy (jdoe0001):

well, in the 1st case, we have a +1, and you distribute inside the parentheses, yes

OpenStudy (anonymous):

so it'd be 2x - 3 > 6?

OpenStudy (jdoe0001):

yes, but solving for "x", "x" would be?

OpenStudy (anonymous):

4.5 and -4.5 ?

OpenStudy (jdoe0001):

well, the 1st case will give you ONLY 9/2 or 4.5 for the second case, you may try to solve for "x" at -1(2x-3) = 6

OpenStudy (jdoe0001):

what would that give you?

OpenStudy (anonymous):

I really dont know..

OpenStudy (jdoe0001):

-1(2x-3) = 6 -2x +3 = 6

OpenStudy (anonymous):

-1.5

OpenStudy (anonymous):

A. x=-1.46 B. x= 6.83 C. x=2 4/7 D. x= -1 3/8 are the ones to choose from

OpenStudy (anonymous):

do u know it?

OpenStudy (jdoe0001):

yes, -3/2 so in the 1st case what happened was $$ |2x-3|>6\\ +1(2x-3)>6\\ 2x-3 > 6 \implies 2x > 9 \implies x > \frac{9}{2}\\ -1(2x-3)>6\\ -2x+3 > 6 \implies -2x > 3\\ \text{the one thing you'd need to keep in mind}\\ \text{for inequalities, is that whenever you}\\ \text{divide or multiply or exponentialize by}\\ \text{a negative value, you have to}\text{ flip } \text{the sign, thus}\\ -2x > 3 \implies x \color{red}{<} -\frac{3}{2} $$